In 5,595 seconds, the number of radioactive nuclei decreases to 1/16 of the number present initially. What is the half-life (in s) of the material? Enter an integer.
Answer #1

1/16 = 4 half life

4 half ife takes = 8854 s

1 half life = 8854/4 = 2213.5 = 2214 s